OpenStudy (anonymous):

SIMPLIFY ~ (x^5)^-8x^4

OpenStudy (anonymous):

\[(x^5)^{-8x^4}=x^{-40x^4}\]

OpenStudy (anonymous):

\[(a^m)^n=a^{mn}\] ^---Rule

OpenStudy (anonymous):

my opinions are fractions.

OpenStudy (anonymous):

\[\frac{1}{x^{40x^4}}\] Is this one of your options?

OpenStudy (anonymous):

no i have 1 over x^160 or 1 over x^36 or x^-36 or i over x^12

OpenStudy (anonymous):

was the question supposed to be \[(x^5)^{-8\times 4}\]?

OpenStudy (anonymous):

If so then the answer is \[\frac{1}{x^{160}}\]

OpenStudy (anonymous):